For each of the following unbalanced reactions, suppose exactly 1.48 g of each reactant is taken. Indicate which reactant is the limiting reagent. Calculate the mass of each product that is expected.

(a) UO2(s) + HF(aq) → UF4(aq) + H2O(l)
limiting reagent

[_]UO2

[_]HF   

UF4 produced g
H2O produced g


(b) NaNO3(aq) + H2SO4(aq) → Na2SO4(aq) + HNO3(aq)
limiting reagent

[_] NaNO3

[_] H2SO4   

Na2SO4 produced g
HNO3 produced g


(c) Zn(s) + HCl(aq) → ZnCl2(aq) + H2(g)
limiting reagent

[_] Zn

[_] HCl   

ZnCl2 produced g
H2 produced g


(d) B(OH)3(s) + CH3OH(l) → B(OCH3)3(s) + H2O(l)
limiting reagent

[_] B(OH)3

[_] CH3OH   

B(OCH3)3 produced g
H2O produced g
Answer #1

(a)

Balanced equation: UO2(s) + 4 HF(aq) → UF4(aq) + 2 H2O(l)

moles UO2 = 1.48 g / 270.02771 g/mol = 0.00548091897 mol

moles HF = 1.48g / 20.006343 g/mol = 0.07397653834 mol

limiting reagent = UO2

mass UF4 = 0.00548091897 mol x 314.022523 g/mol = 1.72113200332

UF4 produced = 1.72 g

H2O produced = 0.197 g

--------------------------------------------------------------------------------------------------------------

(b)

Balanced equation: 2 NaNO3(aq) + H2SO4(aq) → Na2SO4(aq) + 2 HNO3(aq)

moles NaNO3 = 1.48 g / 84.9947 g/mol = 0.01741285044 mol

moles H2SO4 = 1.48g / 98.0785 g/mol = 0.01508995345 mol

limiting reagent = NaNO3

mass Na2SO4 = (1/2) x 0.01741285044 mol x 142.0421 g/mol = 1.23667892174 g

Na2SO4 produced = 1.24 g

HNO3 produced = 1.10 g

--------------------------------------------------------------------------------------------------------------

(c)

Balanced equation: Zn(s) + 2 HCl(aq) → ZnCl2(aq) + H2(g)

moles Zn = 1.48 g / 65.3800 g/mol = 0.02263689201 mol

moles HCl = 1.48g / 36.4609 g/mol = 0.04059142807 mol

limiting reagent = HCl

mass ZnCl2 = 0.02029571403 mol x 136.2860 g/mol = 2.76602168229 g

ZnCl2 produced = 2.77 g

H2 produced = 0.0409 g

--------------------------------------------------------------------------------------------------------------

(d)

Balanced equation: B(OH)3(s) + 3 CH3OH(l) → B(OCH3)3(s) + 3 H2O(l)

moles B(OH)3 = 1.48 g / 61.8330 g/mol = 0.023935439 mol

moles CH3OH = 1.48g / 32.0419 g/mol = 0.04618952059 mol

limiting reagent = CH3OH

mass B(OCH3)3 = 0.01539650686 mol x 103.9128 g/mol = 1.59989413804 g

B(OCH3)3 produced = 1.60 g

H2O produced = 0.832 g

--------------------------------------------------------------------------------------------------------------

Hope this helped you!

Thank You So Much! Please Rate this answer as you wish.("Thumbs Up")